OpenStudy (erinweeks):

Use the compound interest formulas A = Pert and A = P(1+r/n)^nt to solve. Suppose that you have $11,000 to invest. Which investment yields the greater return over 10 years: 6.25% compounded continuously or 6.3% compounded semiannually?

OpenStudy (erinweeks):

A. $11,000 invested at 6.25% compounded continuously over 10 years yields the greater return. B. Both investment plans yield the same return. C. $11,000 invested at 6.3% compounded semiannually over 10 years yields the greater return.

OpenStudy (psymon):

Well for your compounded continuously option, we just plug in the numbers into the A =Pe^(rt) formula. P is your principal, your initial amount, so 11,000. Rate is your percentage written as a decimal, so .0625 and time is 10. This gives us this equation: \[A = 11000e ^{(10)(.0625)}\] For the second equation, r is again written as a decimal, but is .063 this time. Time is still 10 years. N is the amount of times the interest is compounded each year. So semi-annually means twice a year, meaning n = 2. So if we plug the numbers into that formula, we get: \[A = 11000(1 + \frac{ .063 }{ 2 })^{(2)(10)}\] Now you just need to do the calculator work and see which yields the greater return :3

OpenStudy (erinweeks):

whats the "e" in the first equation?

OpenStudy (anonymous):

e is a number kinda like pi.

OpenStudy (anonymous):

2.71...

OpenStudy (psymon):

It stands for exponential function. It's the inverse of the log function. Where it actually comes from is: \[\lim_{n \rightarrow \infty}(1+\frac{ 1 }{ n })^{n}\] A calculus concept, but yeah....2.72blah blah.

OpenStudy (erinweeks):

well for 1. i got 335.637 without doing the e or whatever for 2. i got 1.250^81

OpenStudy (psymon):

You need to calculate the e, though. e is a number that must be calculated. You're taking e to a certain power in the first equation. For #2 you should get something like 20454.

OpenStudy (psymon):

Your calculator should have a button like \[e ^{x} \] on it somewhere

OpenStudy (erinweeks):

ok i see it

OpenStudy (psymon):

Yep. So we need to do e^(.625) and then multiply that result by 11000.

OpenStudy (erinweeks):

for one i got 20550.70553

OpenStudy (psymon):

That sounds right ^_^ Now see if you can do it correctly for the 2nd one.

OpenStudy (erinweeks):

i got 21887.67749

OpenStudy (psymon):

Hmm...Wonder why you're getting that. Probably the order of operations make it harder on youf or your calculator.

OpenStudy (psymon):

So let's do this piece by piece then, starting with inside of the parenthesis. So if I only do: \[1+\frac{ .063 }{ 2 }\], you should get 1.0315, correct?

OpenStudy (erinweeks):

yes!

OpenStudy (psymon):

Awesome. Now take that to the 20th power.

OpenStudy (erinweeks):

1.859450605

OpenStudy (psymon):

Multiply by 11000.

OpenStudy (erinweeks):

20453.95666

OpenStudy (psymon):

Much better :P So you can see which one is larger now :3

OpenStudy (erinweeks):

lol sorrry :p and yes the first ((:

OpenStudy (psymon):

Yeah, so there ya go :3

OpenStudy (erinweeks):

thankk yaa once again !

OpenStudy (psymon):

yep yep :3